Your boss asked you to evaluate a project with an infinite life. Sales and costs project to $1,000 and $500 per year, respectively. (Assume sales and costs occur at the end of the year [i.e., profit of $500 at the end of year one]). There is no depreciation and the tax rate is 30 percent. The real required rate of return is 10 percent.The inflation rate is 4 percent and is expected to be 4 percent forever. Sales and costs will increase at the rate of inflation. If the project costs $3,000, what is the NPV?

A.$500.00

B. $1,629.62

C. $365.38

D. $472.22

E. I choose not to answer

I see that some have answered this question and taken A. I am wondering, why are we not changing the real rate to nominal as our teacher says we always need to assume the numbers are nominal is it does not say anything. Because If I change the rate to nominal my answer is C. Is it some rule that if inflation grows constant we are not suppose to change the rate?

Thank you for the explanation in advance :)
Answer #1

Year Cash flow 0 -3000 thereafter 350 =500*0.7 So Npv using real return of 10% 10 npv 500 11 12 Nominal return 0.144 =1.1*1.0b